# Perceptions of Linguistic Uncertainty by Language Models and Humans

### v1

@@ B.4 Prompts
Figure 11: Structure of a demonstration (or exemplar). The placeholders [[speaker]] are replaced by gendered names, [[uncertainty]] is replaced by the uncertainty expressions, and [[statement]] is replaced with the corresponding statements.
During the course of our experiments, we carried experiments with varying assumptions: non-verifiable setup assessed models (and humans) perceptions in the absence of strong prior knowledge about the statements, whereas the verifiable setup focused on the evaluation of the same perceptions when knowledge was present. We used two different sets of exemplars in our experiments to reflect these differences. The set of exemplars used in the non-verifiable setting are defined in terms of the following:
@@
As for the verifiable exemplars, we experimented with various orderings of the statements and found minimal differences in the obtained distribution of numerical responses. As a result, we report the results using a false statement as the first example and a true statement as the second example in the prompt. These were associated with the following speaker names and uncertainty expressions:
@@
B.5 Non-Verifiable Statements
In this section, we list a random selection of 20 statements (5 per each statement type). These statements are incorporated in the prompts listed in the main paper and the placeholders [[they]] and [[their]] are replaced by pronouns matching the gender of the statement speaker’s name.
Forecasting of future events.
Verbal probabilities are often used to communicate uncertainty about future events.
  1. 1.
    [[they]] will buy a new watch this Thanksgiving weekend.
  2. 2.
    [[they]] will be offered a promotion this fall.
  3. 3.
    the company will have another round of lay-offs by mid July.
  4. 4.
    there will be vegetarian options at the barbecue.
  5. 5.
    [[they]] will visit New York over winter break.
Imperfect knowledge.
Verbal probabilities can also be used to communicate uncertainty imprecise information about events or outcomes.
  1. 1.
    the restaurant near [[their]] apartment accepts reservations.
  2. 2.
    the new museum is offering complimentary admission.
  3. 3.
    there is a yoga studio within 2 miles of [[their]] workplace.
  4. 4.
    there are more than eighty students in the auditorium right now.
  5. 5.
    the temperature in the office is at least 72 degrees Fahrenheit.
Possession.
Alternatively, verbal probabilities can be used to convey uncertainty about acquaintances, be it in terms of the objects they own or in terms of their preferences.
  1. 1.
    [[their]] boss owns a blue car.
  2. 2.
    [[their]] friend has a leather jacket.
  3. 3.
    [[their]] cousin has a vegetable garden.
  4. 4.
    [[their]] classmate owns a guitar.
  5. 5.
    [[their]] boss has a stereo amplifier.
Preference.
  1. 1.
    [[their]] cousin prefers spinach over broccoli.
  2. 2.
    [[their]] boss prefers coffee over tea.
  3. 3.
    [[their]] friend prefers running over cycling.
  4. 4.
    [[their]] neighbor prefers the beach over the mountains.
  5. 5.
    [[their]] coworker prefers reading books over watching movies.
B.6 Verifiable Statements
In this section, we list a random selection of 9 true statements (3 per each topic) and their false counterparts. These statements are incorporated then used as part of the test examples the prompts listed in Section B.4.
Geography.
One of the topics of the experiment involves geography, as well as knowledge about landmarks and monuments. These statements were curated from a set of easy trivia questions provided by The Question Company (as described in Section 4). For each question-answer pair in the trivia dataset, we create both a true and a false statement using the correct and one incorrect answer choice, respectively.
Given our interest in attesting the knowledge effect in the models’ and humans’ perceptions of linguistic uncertainty, we purposely decided to use easy trivia questions as the basis for our facts (as opposed to more difficult ones), since this subset constitutes a good proxy for facts that LLMs and humans may have strong prior beliefs about.
  1. 1.
    Great Britain directly borders 0 countries.
  2. 2.
    the Colosseum, a famous landmark in Rome, was originally built as an Amphitheatre.
  3. 3.
    New York is known as the Big Apple.
  4. 4.
    Great Britain directly borders 2 countries.
  5. 5.
    the Colosseum, a famous landmark in Rome, was originally built as an Cathedral.
  6. 6.
    New York is known as the Big Orange.
History of Art.
One of the topics of the experiment involves history of arts. For each fact we include both a true and one false variation of that fact.
  1. 1.
    the Mona Lisa is a famous painting by Leonardo da Vinci.
  2. 2.
    the Scream is the best known painting by Edvard Munch.
  3. 3.
    Andy Warhol became a famous artist in the 1960s for painting soup cans and soap boxes.
  4. 4.
    the Mona Lisa is a famous painting by Tintoretto.
  5. 5.
    the Scream is the best known painting by Jackson Pollock.
  6. 6.
    Frida Kahlo became a famous artist in the 1960s for painting soup cans and soap boxes.
Science.
These include facts concerning chemistry, biology, and astronomy. For each fact we include both a true and one false variation of that fact.
  1. 1.
    water’s chemical formula is H2O.
  2. 2.
    pH is a measure of the acidity or basicity of a substance.
  3. 3.
    the nearest planet to the sun is Mercury.
  4. 4.
    carbon monoxide’s chemical formula is H2O.
  5. 5.
    oG is a measure of the acidity or basicity of a substance.
  6. 6.
    the nearest planet to the sun is Mars.
B.7 Language Models
Throughout our paper, we use OpenAI to obtain the results for ChatGPT, GPT-4, and GPT-4o; Google’s Vertex AI APIs to obtain results for Gemini, TogetherAI [^12] https://www.together.ai/ to run LLama3 (70B), Mixtral 8x7B, and Mixtral 8x22B. We run LLama3 (8B) OLMo (7B) and Gemma (2B) locally on a single GPU 8 RTX A6000 (48 GB).
During the paper, we shorten the name of the studied models for simplicity. All our experiments consider the instruction-tuned or RLHF version of the mentioned models. All experiments were conducted from April through June. For reproducibility, we list below the mapping from model name to exact version of the model used:
@@ B.7 Language Models
Appendix C Extracting LLMs Numerical Responses
In this section, we elaborate on the methodologies to extract the numerical responses. There are many different level of information that we can use to characterize numerical responses from LLMs. When estimating empirical distributions using point estimates, it suffices to use greedy decoding, since it represents the best guess that a model would give on average. However, in some cases, one may be interested in more faithful estimates of the empirical probability distribution for each uncertainty expression. To that end, one can take advantage of the information available in different platforms, which we describe next.
C.1 Full Next-Token Probability Information
By definition, our task elicits a numerical response from LLMs, which resembles the setup in verbalized confidence (Tian et al., 2023). The adoption of single digit tokenization (Singh and Strouse, 2024) by autoregressive models (e.g., Gemma (2B), LLama3 (70B), and OLMo (7B)) creates some challenges in the computation of numerical responses non-trivial for autoregressive models. In practice, due to the left-to-right nature of LLMs, single digit tokenization implies that the probability of a number between $[0,9]$ is always greater or equal to the probability of any number in $[10,100]$. To circumvent this problem, we report the corrected probability during our experiments as follows:
$p_{\mathrm{model}}(y_{t}=i|x)-\sum_{j=0}^{9}p_{\mathrm{model}}(y_{t}=i,y_{t+1}=j|x),$
where $x$ is a prompt and $j\in[0,9]$. Intuitively, this means that we are computing the probability of $i\in[0,9]$ and no other number following it. The details of what a number is change with tokenizer implementation. Unlike traditional greedy decoding, we condition the selection of the arg-max prediction to the set of strings representing the numbers between $[0,100]$ (followed by no other number).
C.2 Partial Next-Token Probability Information
@@
In order to work, this method requires two properties to be satisfied: (1) numbers between 0 and 100 were encoded with unique tokens (i.e., there are 101 unique integers that represent each individual token), and (2) exponentiating the log probabilities returned by the black-box API must lead to a valid probability distribution (i.e., numbers obtained for different prompts will be comparable to one another). For OpenAI models, the first requirement is satisfied.
@@
Selecting the greedy prediction: Unlike traditional greedy decoding, we condition the selection of the arg-max prediction over numbers the top-k (k=20 for OpenAI). That is, we select the most likely number that is present in the top-20 predicted tokens.
Estimating conditional distributions using probabilistic decoding: In Section 5.4, we use the information available in the top-20 tokens made available by OpenAI models. Like before we bin the predictions into 21 bins, defined from 0 to 100 in increments of 5. To ensure that we have a valid probability definition, we consider an additional bin (“-1”) that accumulates the probability of not having being a number in the top-20.
@@
Figure 12: Differences in the probability mass as determined by OpenAI models on the top-20 tokens. We report these values across all statements (n=840) in the verifiable and non-verifiable settings. We observe that numbers account for the majority of probability mass in ChatGPT and GPT-4. Upon analysis we found lower probability mass assigned by GPT-4 to be correlated with the appearance of the words “Given”, “The”, and “And”.
In Section 4.2 we briefly mention the use of different methodologies to extract numerical responses from LLMs. The following describes the list of methodologies used for each LLM:
@@
  • Greedy Sampling: Gemini, LLama3 (70B), Mixtral 8x7B, Mixtral 8x22B. We opt for using greedy sampling as opposed to standard sampling due to budget constraints. Given the nature of our experiments, faithfully estimating the empirical distributions over the 100 numbers would require hundreds or thousands of calls. These calls are time-consuming and costly. We believe that using decoding is still representative of how a model would behave in most cases.
@@
  • Full next-token probability distribution: LLama3 (8B), Gemma (2B), OLMo (7B). We found these models to be particularly brittle to the prompts.
@@
  • Next-token probability distribution: ChatGPT, GPT-4, GPT-4o. As of June 2024, OpenAI models only provide access to the next-token probabilities of the top-20 tokens. During the experiments in Section 5.4, we collect the information about the top 20 numbers
Appendix D Additional Results
In this section, we report additional results, including visualization of the empirical distributions for models and humans in Section D.1, additional measurements of similarity between non-verifiable and verifiable distributions in Section D.2, metrics discriminated by uncertainty expression (whenever applicable) in Sections D.3 and D.4.
D.1 Histograms
(a) Humans
(b) Random Baseline
(c) Gemma-1.1-2b-it
(d) OLMo-7B-instruct
(e) gemini-pro
(f) GPT-4o
(g) ChatGPT
(h) GPT-4
(i) LLama3 (8B)
(j) LLama3 (70B)
(k) Mixtral 8x7B
(l) Mixtral 8x22B
Figure 13: Empirical distributions of numerical probabilities per uncertainty expression in the non-verifiable setting. For each uncertainty expression (row), the empirical distribution is computed based on n=60 datapoints (for LLMs) and n=188 (for humans). We observe that larger, newer LLMs match the mode of human behavior.
D.2 Summary Metrics
For a more complete understanding of the differences among the distributions, we report distance metrics in Table 4.
  • Proportional Agreement: proposed in Section 4.3, measures the overall agreement between an agent’s and a reference (population) distribution. We use the results of the human studies in the non-verifiable setting as our reference distribution throughout the whole paper.
  • Mean Absolute Error: proposed in Section 4.3, measures the average agreement across uncertainty expressions between a agent’s distribution and a reference (population) distribution. In this case, we also use the human results from the non-verifiable setting as our reference distribution throughout the paper.
  • Wasserstein Distance: computed using scipy.stats.wasserstein_distance, measures the distance between two conditional distributions.

### v2

@@ C.4 Prompts
Figure 11: Structure of a demonstration (or exemplar), as defined in the 2-shot prompt. The placeholders [[speaker]], [[uncertainty]], and [[statement]] are replaced by gendered names, uncertainty expressions, and statements, respectively.
During the course of our experiments, we carried experiments with varying assumptions: non-verifiable setup assessed models (and humans) perceptions in the absence of strong prior knowledge about the statements, whereas the verifiable setup focused on the evaluation of the same perceptions when knowledge was present. We used two different sets of exemplars (or demonstrations) in our experiments to reflect these differences. The set of exemplars used in the non-verifiable setting are defined in terms of the following:
@@
Conversely, for the set of exemplars used in the verifiable setting, we carried a preliminary experiment to determine whether the truthness/falsehood of the examples and their ordering had a significant impact in models’ performance. We used the two pairs of statements (“only some metals can conduct electricity,” “all metals can conduct electricity”) and (“the Sun orbits around the planet Earth,” “the planet Earth orbits around the Sun”) in the preliminary experiments. We found negligible differences in the obtained distribution of numerical responses (as emphasized in Tables 7 and 8). As a result, we report the results using a false statement as the first example and a true statement as the second example in the prompt (denoted FT). These were associated with the following speaker names and uncertainty expressions:
@@
Table 7: Average distributional difference in LLMs conditional distributions when estimated based on different true/false configurations of the examples in the verifiable prompt. We report the Wasserstein distance of the estimated conditional distributions averaged over the empirical distributions obtained using the other configurations and greedy decoding algorithm (see Table 8 for results using the probabilistic decoding). On average, we observe that changing the veracity of the examples in the prompt has minimal impact on the final distributions. Moreover, we explicitly avoid using the configurations FF or TT to avoid majority label bias (Zhao et al., 2021).
ChatGPT GPT-4 GPT-4o
FF 2.33 1.31 0.92
FT 3.20 1.25 0.85
TF 2.93 1.32 0.88
TT 2.31 1.38 0.90
Table 8: Average distributional difference in LLMs conditional distributions when estimated based on different true/false configurations of the examples in the verifiable prompt. On average, we observe that changing the veracity of the examples has minimal impact on the final distributions.
ChatGPT GPT-4 GPT-4o
FF 1.91 1.53 1.74
FT 3.01 1.36 1.36
TF 2.77 1.33 1.81
TT 1.91 1.48 1.39
C.5 Language Models
Our evaluation concerns the study of instruction-tuned LLMs, some of which are accessible through black-box APIs and others through the use of the HuggingFace Python package. We use OpenAI to obtain the results for ChatGPT, GPT-4, and GPT-4o; Google’s Vertex AI API to obtain results for Gemini, and TogetherAI [^12] https://www.together.ai/ to run LLama3 (70B), Mixtral 8x7B, Mixtral 8x22B, and DBRX. We run LLama3 (8B) OLMo (7B) and Gemma (2B) locally on a single GPU 8 RTX A6000 (48 GB). All experiments were conducted from April through June 2024.
For simplicity, we have abbreviated the name of the evaluated models, removing information about the size and version of the model. For reproducibility, we list below the mapping from model name to exact version of the model used:
@@ C.5 Language Models
Appendix D Extracting LLMs Numerical Responses
In this section, we outline the methodologies used to extract the numerical responses from auto-regressive LLMs. In an ideal world, given an uncertainty expression $u$, the model’s conditional distribution $\hat{P}_{\mathrm{model}}(k|u)$, would be fully observable. However, information about $\hat{P}_{\mathrm{model}}$ is seldom available, for models are either (1) served through opaque closed-source APIs or (2) too large, often requiring large amounts of compute to estimate the full distribution. To circumvent such limitations, one idea is to empirically estimate such probabilities using sampling-based approaches (e.g., self-consistency (Wang et al., 2023a)) or using greedy decoded sequences over multiple examples. In the ensuing sections, we describe the three different strategies considered in our work to estimate the LLMs’ empirical distributions.
D.1 Method 1: Full Probability Approach
The “full” methodology requires access to the next-token probability distribution of an auto-regressive LLM and, as a result, is currently applicable to open source models. Because we use the models’ probability to compute the probability of producing any integer between 0 and 100, it is also our most time-consuming approach [^13] Running LLama3 (70B) across 5x GPU 8 RTX A6000 (48 GB) for 900 examples required approximately 9 full days. For that reason, we only apply this strategy to a few LLMs., requiring 101 model calls to obtain the full probability.
Methodology. An important aspect to account for when using LLMs to estimate the probability distribution over the set of integers ranging from 0 to 100 is the LLM’s tokenizer. Specifically, models, such as Gemma (2B), LLama3 (70B), and OLMo (7B), use single-digit tokenization (Singh and Strouse, 2024), which implies that the textual representation of 100 (represented as “100”) is tokenized into at least three individual tokens (i.e., “1”, “0”, “0”) and not in a single token (e.g., “100”). In terms of probabilities, this is a challenge, since many of these LLMs generate digits in a left-to-right fashion and, by design, the probabilities of any number in $[10,100]$ are always less than or equal to the probability of the first constituting digit (e.g., probability of “10” subsumes the probability of “1” since in order to generate the string “10” the model needs to generate the string “1” first). Since we are interested in knowing the true independent probability assigned to every integer between 0 and 100, we need to adjust the LLMs’ probability. Therefore, instead of reporting the probability that a number $i\in[0,100]$ occurs, we compute the probability that $i$ occurs and is not followed by a number $j\in[0,9]$:
$p(x_{t}=i|x_{<t})-\sum_{j=0}^{9}p(x_{t}=i,x_{t+1}=j|x_{<t}),$
where $p$ represents the LLM’s next-token probability distribution. In sum, given a prompt parameterized with a speaker $s$, uncertainty expression $u$, statement $e$, and a prompt that combines these parameters, denoted $\text{prompt}(s,u,e)$ we use the expression above to obtain the full probability distribution for all $k\in 0,...,100$. We denote this corrected probability distribution as $p(k|\text{prompt}(s,u,e))$ and we refer to the probability mass that falls outside the set of strings $k\in 0,...,100$ as $\bar{p}$.
Constructing the Greedy Histogram. Unlike sampling-based greedy decoding algorithms, we restrict the selection of the arg-max to the set of strings representing the numbers between $[0,100]$, as opposed to sampling a series of tokens using temperature=0. In other words, we constrain the greedy decoding to be any of the sequences in {“0”, “1”, …, “100”}, regardless of how little probability [^14] We found this approach to be particular sensitive to the instruction format used. For example, in earlier iterations of this work we used Llama-2 and Gemma (7B) but found them to be particular sensitive to the whitespaces provides in the prompt. is assigned to any of these numerical sequences. For every triplet $(u,s,e)$, we obtain the the arg-max prediction and then assign it to a bin $0,5,10,...,100$.
Constructing the Probabilistic Histogram. The probabilistic histogram benefits from the probability information that was computed previously for a specific triplet $(u,s,e)$. In particular, for a specific example we accumulate all probability values in the corresponding bins $0,5,...,100$. The remaining probability mass $\bar{p}$ is assigned to a default bin “-1”. In other words, the bin “-1” will accumulate the probability of a number in [0, 100] not following the specified prompt. While we could have normalized the probabilities of $0,5,...,100$ to sum to 1, we decided to add a “-1” bin to allow for a fair comparison with the top-k approach, where only part of the probability distribution (e.g., top-20 next probability distribution values are revealed), as is the case for the OpenAI models. After accumulating the probability over all $(u,s,e)$ triplets, we normalize the histogram by dividing by the number of triplets.
D.2 Method 2: Top-k approach
This method leverages information about the top-k values of the next-token probability information. Applicable to OpenAI models, namely ChatGPT, GPT-4, and GPT-4o, this approach allow us to obtain richer probability information with a single API call.
@@ D.2 Method 2: Top-k approach
Methodology. This method requires two properties to be satisfied: (1) numbers between 0 and 100 must be encoded with one single token each (i.e., each integer is represented with a single token), and (2) exponentiating the log probabilities returned by the API must lead to a valid probability distribution (i.e., numbers obtained for different prompts will be comparable to one another). We validate that the first requirement is satisfied by OpenAI models.
@@
Constructing the Greedy Histogram. Unlike traditional greedy decoding, we condition the selection of the arg-max prediction over numbers the top-k (k=20 for OpenAI). That is, we select the most likely number that is present in the top-20 predicted tokens. If no number is present in the top-20 tokens, we assign a default value of ‘-1’.
Constructing the Probabilistic Histogram. Like the “full” probability approach, we construct the probabilistic histogram by accumulating the probabilistic information that we gather with each inference call: For a given triplet $(u,s,e)$, we make an API call and obtain probability about the next 20 tokens. If any of these strings represents a number between 0 and 100, we exponentiate it to obtain a probability value, and accumulate the probability in the corresponding bin. Any remaining probability that is not assigned to a number in the top-20 is accumulated in the ‘-1’ bin. After accumulating the probability over all $(u,s,e)$ triplets, we normalize the histogram by dividing by the number of triplets.
@@
Figure 12: Differences in the probability mass as determined by OpenAI models on the top-20 tokens. We report these values across all statements (n=840) in the verifiable and non-verifiable settings. We observe that numbers account for the majority of probability mass in ChatGPT and GPT-4. Upon analysis we found lower probability mass assigned by GPT-4 to be correlated with the appearance of the words “Given,” “The,” and “And”.
D.3 Method 3: Sampling-based approach
Some models are provided through black-box APIs with no access to next-token probability information. Ideally, one could estimate the next-token probability information by continuously sample, but doing so would be too costly. In the main paper, we only run sampling based approaches using a single sample.
Constructing the Greedy Histogram. This approach differs from the previous two in that it is unconstrained sampling-based approach, which means that the next immediate token may not be a number. Moreover, it is also agnostic to the tokenization. It many cases, the model may generate some text before it actually produces the answer “The answer is 100” or “**Sure, here is the answer: 55”. To accommodate for such scenarios, we consider the arg-max prediction to be the first number between 00 and $100$ to be mentioned in the model’s response. Otherwise we assume the arg-max prediction to be ‘-1’.
Constructing the Probabilistic Histogram. Due to budget constraints, we could not reliably estimate a probabilistic histogram via the sampling-based approach, as this would require thousands of requests. Future work could use strategies like self-consistency to estimate a probabilistic histogram.
D.4 Which methodology applied to which model?
The following describes the list of methodologies used for each LLM:
@@ D.4 Which methodology applied to which model?
  • Greedy Sampling: Gemini, LLama3 (70B), Mixtral 8x7B, Mixtral 8x22B. We opt for using greedy sampling (max_tokens=200, temperature=0) as opposed to standard sampling due to budget constraints. Given the nature of our experiments, faithfully estimating the empirical distributions over the 101 numbers would require hundreds or even thousands of calls. These calls are time-consuming and costly. We believe that using greedy decoding is still representative of how a model would behave in the majority of the cases.
@@
  • Full next-token probability distribution: LLama3 (8B), Gemma (2B), OLMo (7B). We found these models to be particularly brittle to prompting.
@@
  • Next-token probability distribution: ChatGPT, GPT-4, GPT-4o. As of June 2024, OpenAI models only provide access to the next-token probabilities of the top-20 tokens. In the experiments in Section 5.4, we collect the information about the top 20 numbers.
Appendix E Additional Results
In this section, we report the following additional results:
  • In Section E.1 we report the mean and 95% confidence estimates for the proportional agreement (PA) and mean average error (MAE) metrics. Since the PA prioritizes mode-matching behavior, we also include a distribution-matching metric based on Wasserstein distance.
  • In Section E.2 we analyze differences in model responses between true and false statements on a pairwise basis, pairing statements by the original question they were based on.
  • In Section E.3 we comment on the variability of the distributions obtained when using probabilistic decoding vs simple greedy decoding.
  • In Section E.4 we provide visualizations of the histogram distributions for the non-verifiable and verifiable settings.
E.1 Mode- and Distribution-Matching Metrics
Tables 9 and 10 show the mean and 95% confidence intervals for the PA and MAE metrics. For completeness, we also include the PA score and MAE metrics per expression in Sections E.1.1 and E.1.2.
PA (95% CI) % PA MAE (95% CI)
Human Mode 27.6
Human Individual $17.6_{(17.2,18.0)}$ 63.8 $8.91_{(8.08,9.17)}$
[…]
GPT-4 $24.4_{(24.2,24.6)}$ 88.4 $4.64_{(4.53,4.74)}$
GPT-4o $18.9_{(18.4,19.5)}$ 68.5 $5.58_{(5.34,5.79)}$
Gemini $25.4_{(25.1,25.6)}$ 92.0 $4.09_{(3.92,4.23)}$
LLama3 (8B) $17.8_{(17.4,18.2)}$ 64.5 ${11.99}_{(10.92,13.13)}$
LLama3 (70B) $23.6_{(23.4,23.8)}$ 85.5 ${5.56}_{(5.34,5.80)}$
Mixtral 8x7B $21.8_{(21.4,22.0)}$ 79.0 ${5.88}_{(5.44,6.34)}$
Mixtral 8x22B $21.8_{(21.6,22.0)}$ 79.0 ${7.20}_{(6.56,7.81)}$
OLMo (7B) $11.1_{(10.7,11.5)}$ 40.2 ${21.41}_{(20.17,22.41)}$
Gemma (2B) $8.1_{(7.7,8.5)}$ 29.3 ${20.17}_{(19.11,21.17)}$
PA (95% CI) $\Delta$ PA MAE (95% CI) $\Delta$ MAE
Human Mode 27.6
Human Individual $16.7_{(16.3,17.1)}$0.9 $9.35_{(8.23,9.50)}$ 0.44
[…]
GPT-4 ${22.1}_{(21.7,22.4)}$2.3 ${3.84}_{(3.03,4.45)}$0.80
GPT-4o ${15.2}_{(14.5,15.9)}$3.7 ${7.05}_{(6.45,7.62)}$ 1.47
Gemini ${21.3}_{(20.7,21.8)}$4.1 ${7.23}_{(6.04,8.49)}$ 3.14
LLama3 (8B) ${10.1}_{(9.5,10.7)}$7.7 ${16.59}_{(15.10,18.18)}$ 4.60
LLama3 (70B) ${18.9}_{(18.1,19.6)}$4.7 ${13.73}_{(11.92,15.64)}$ 8.17
Mixtral 8x7B ${15.2}_{(14.5,15.9)}$6.6 ${12.23}_{(10.37,14.18)}$ 6.35
Mixtral 8x22B ${18.6}_{(18.3,19.0)}$3.2 ${9.78}_{(8.51,11.08)}$ 2.58
OLMo (7B) ${7.6}_{(7.2,8.0)}$3.5 ${33.66}_{(31.82,35.11)}$ 12.25
Gemma (2B) ${5.3}_{(5.0,5.6)}$2.8 ${25.07}_{(23.45,26.66)}$ 4.9